# Thunder-NUBench: A Benchmark for LLMs’ Sentence-Level Negation Understanding

### v3

@@ 8 Paraphrased sentence:
Qwen2.5- 3B baseline zero-shot 0.542 67.69 26.02 6.29 24.68 30.19 55.10 42.26
1-shot 0.522 67.63 25.53 6.84 24.36 28.25 55.78 38.06
5-shot 0.481 70.84 23.89 5.27 25.00 28.57 50.00 37.74
10-shot 0.437 73.14 22.14 4.72 23.40 25.32 45.58 38.06
sft zero-shot 0.276 85.34 13.22 1.44 14.74 20.78 30.95 30.97
Qwen2.5- 3B-Instruct baseline zero-shot 0.466 68.03 28.57 3.40 19.23 28.25 42.86 40.97
1-shot 0.393 67.68 28.28 4.04 16.67 20.45 42.52 30.65
5-shot 0.355 77.01 21.88 1.12 19.55 22.40 39.46 31.94
10-shot 0.351 77.83 21.04 1.13 21.15 23.05 35.71 32.90
sft zero-shot 0.300 85.19 13.49 1.32 15.71 22.40 33.33 34.19
Qwen2.5- 7B baseline zero-shot 0.520 70.73 24.54 4.73 24.68 31.17 53.06 43.55
1-shot 0.534 71.47 25.56 2.97 29.17 31.17 56.12 41.61
5-shot 0.492 71.13 25.81 3.06 24.36 28.25 51.70 40.65
10-shot 0.447 74.07 23.98 1.95 24.04 24.68 45.58 42.58
sft zero-shot 0.240 87.75 11.59 0.66 15.06 15.91 22.45 33.23
Qwen2.5- 7B-Instruct baseline zero-shot 0.458 65.16 31.54 3.29 18.27 29.22 39.80 36.13
1-shot 0.438 69.02 30.07 0.91 22.12 23.70 43.88 35.48
5-shot 0.378 71.85 27.94 0.21 19.55 19.16 37.41 36.13
10-shot 0.335 73.76 25.77 0.47 19.23 18.18 30.61 34.19
sft zero-shot 0.282 88.76 10.96 0.28 16.03 19.48 31.29 36.77
Table 21: Error rates, incorrect choice distributions, and local negation confusion rates for the Qwen family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
@@
Qwen2.5- 3B baseline zero-shot 0.500 0 38.99 19.49 41.52 23.08 19.16 26.87 11.61
1-shot 0.463 0 45.89 27.23 26.88 25.00 21.43 26.53 14.84
5-shot 0.431 0 44.75 25.60 29.65 24.04 19.81 23.13 12.58
10-shot 0.439 0 45.85 23.10 31.05 25.00 21.10 25.17 11.94
sft zero-shot 0.373 2 51.07 13.25 35.68 17.31 15.58 26.19 19.35
Qwen2.5- 3B-Instruct baseline zero-shot 0.380 0 49.27 40.29 10.44 20.19 15.91 24.15 17.10
1-shot 0.419 0 53.22 32.39 14.39 26.28 19.16 28.23 18.39
5-shot 0.398 2 54.40 31.60 14.00 27.24 19.16 29.25 13.55
10-shot 0.388 0 56.44 29.24 14.31 26.60 20.13 30.61 13.23
sft zero-shot 0.274 1 59.13 34.49 6.38 12.18 13.31 21.43 20.00
Qwen2.5- 7B baseline zero-shot 0.377 0 68.21 15.79 16.00 28.53 22.73 36.39 18.71
1-shot 0.392 0 70.65 15.59 13.77 33.97 26.30 38.10 16.13
5-shot 0.377 0 72.00 15.16 12.84 33.65 23.05 38.78 16.77
10-shot 0.372 0 73.13 14.93 11.94 32.69 23.05 40.48 16.45
sft zero-shot 0.344 0 82.03 13.13 4.84 30.77 24.35 43.54 18.39
Qwen2.5-7B- Instruct baseline zero-shot 0.324 0 62.84 29.34 7.82 23.72 22.73 23.47 14.19
1-shot 0.363 0 61.79 31.66 6.55 30.45 21.10 27.21 13.87
5-shot 0.368 0 66.38 26.51 7.11 31.41 19.81 34.69 15.16
10-shot 0.390 0 67.07 25.81 7.11 34.62 20.78 35.03 17.74
sft zero-shot 0.207 0 78.54 17.62 3.83 18.59 20.13 17.69 10.65
Table 22: Error rates, incorrect choice distributions, and local negation confusion rates for the Qwen family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting.
Error
Rate
Answer Format Wrong Incorrect Choice Distribution Local Negation Confusion Rate
(1-acc)
@@
gpt-4o- mini baseline zero-shot 0.357 0 34.67 65.33 0 17.31 8.44 13.27 11.94
1-shot 0.317 0 40.00 59.75 0.25 19.87 7.47 13.95 10.97
5-shot 0.316 0 38.35 61.15 0.50 19.23 7.14 11.90 11.61
10-shot 0.298 0 42.02 57.98 0 19.23 8.12 12.93 11.29
gpt-4.1- mini baseline zero-shot 0.203 0 51.95 48.05 0 14.10 6.49 8.16 14.52
1-shot 0.184 0 64.66 35.34 0 16.35 6.17 10.20 16.13
5-shot 0.156 0 65.48 34.52 0 13.46 6.17 7.14 15.16
10-shot 0.153 0 64.77 35.23 0 14.74 5.19 4.76 15.81
claude-3-5- haiku-latest baseline zero-shot 0.278 0 67.81 30.48 1.71 22.76 14.29 15.99 24.52
1-shot 0.477 434 83.93 15.48 0.60 20.19 6.49 10.54 8.71
5-shot 0.527 574 95.56 4.44 0 13.78 3.57 5.44 5.16
10-shot 0.368 324 94.29 5.00 0.71 20.83 6.17 5.44 10.32

### v4

@@ 8 Paraphrased sentence:
Qwen3-0.6B-Base baseline zero-shot 0.577 71.53 22.97 5.50 29.81 35.39 54.42 50.97
1-shot 0.502 73.46 21.80 4.74 29.17 30.84 50.34 42.26
5-shot 0.426 77.28 19.18 3.54 23.40 23.70 44.22 44.84
10-shot 0.401 79.01 18.61 2.38 22.12 23.70 42.18 42.90
after SFT zero-shot 0.277 83.67 15.47 0.86 17.31 17.53 28.23 32.58
Qwen3-0.6B baseline zero-shot 0.587 77.70 17.57 4.73 27.24 41.56 62.93 57.10
1-shot 0.590 76.48 18.82 4.70 30.77 38.64 64.29 53.23
5-shot 0.517 78.53 17.18 4.29 25.00 31.82 57.82 53.55
10-shot 0.471 81.31 15.15 3.54 25.64 27.92 52.38 52.58
after SFT zero-shot 0.321 84.44 14.07 1.48 17.63 20.13 35.03 39.35
Qwen3-1.7B-Base baseline zero-shot 0.538 71.24 22.71 6.05 29.81 33.12 58.50 37.42
1-shot 0.479 74.17 22.85 2.98 28.21 30.52 53.06 35.48
5-shot 0.417 77.38 20.53 2.09 23.40 26.30 41.84 41.94
10-shot 0.401 80.40 18.22 1.39 23.72 24.68 41.84 42.90
after SFT zero-shot 0.276 85.63 12.93 1.44 17.95 24.03 31.97 23.87
Qwen3-1.7B baseline zero-shot 0.581 67.76 25.27 6.97 27.88 37.66 63.27 34.52
1-shot 0.570 71.35 24.34 4.31 29.49 33.77 61.22 44.19
5-shot 0.504 76.54 21.10 2.36 26.60 26.95 58.16 48.06
10-shot 0.485 77.45 20.42 2.12 27.56 25.32 54.42 48.39
after SFT zero-shot 0.358 82.71 15.74 1.55 18.59 26.30 39.12 38.39
Qwen3-4B-Base baseline zero-shot 0.515 75.50 21.57 2.93 27.56 30.84 55.10 47.42
1-shot 0.494 75.12 22.15 2.73 29.49 29.55 52.72 41.94
5-shot 0.449 74.20 22.79 3.00 25.00 27.27 44.22 41.29
10-shot 0.440 76.22 21.80 1.98 25.64 25.97 45.58 41.61
after SFT zero-shot 0.271 86.26 12.87 0.88 16.67 19.48 31.63 29.03
Qwen3-4B baseline zero-shot 0.562 73.73 20.34 5.93 28.21 38.96 65.31 39.35
1-shot 0.495 74.68 20.19 5.13 28.85 29.87 57.82 36.77
5-shot 0.431 78.12 19.12 2.76 23.08 25.65 51.70 39.35
10-shot 0.398 79.08 18.53 2.39 21.79 25.65 46.60 36.45
after SFT zero-shot 0.313 83.76 14.47 1.78 16.67 21.10 28.91 41.29
Qwen3-8B-Base baseline zero-shot 0.545 74.09 21.69 4.22 29.17 33.44 61.22 43.55
1-shot 0.489 73.42 23.18 3.40 27.56 28.25 55.44 37.74
5-shot 0.436 73.09 23.27 3.64 22.44 24.35 44.90 40.32
10-shot 0.412 76.35 21.15 2.50 22.76 25.65 40.48 41.29
after SFT zero-shot 0.286 86.43 12.47 1.11 16.35 20.45 34.01 31.61
Qwen3-8B baseline zero-shot 0.527 71.84 23.04 5.12 26.60 35.06 61.22 34.19
1-shot 0.484 75.41 20.49 4.10 31.73 30.19 57.14 32.26
5-shot 0.442 80.43 17.59 1.97 27.56 30.52 49.32 39.68
10-shot 0.408 81.36 16.70 1.94 25.64 28.57 45.24 38.06
after SFT zero-shot 0.260 83.84 14.33 1.83 13.78 17.21 27.55 31.61
Qwen3-14B-Base baseline zero-shot 0.493 70.69 23.67 5.64 25.00 31.17 46.60 41.29
1-shot 0.443 74.37 22.58 3.05 25.64 27.27 46.26 37.10
5-shot 0.369 76.99 20.65 2.37 20.19 22.08 35.37 39.68
10-shot 0.335 77.07 21.75 1.18 20.51 19.81 28.91 37.42
Qwen3-14B baseline zero-shot 0.512 69.35 23.37 7.28 25.64 33.12 56.46 32.26
1-shot 0.470 75.89 18.89 5.23 31.09 30.84 55.78 30.32
5-shot 0.397 79.60 18.60 1.80 23.72 25.65 44.90 36.45
10-shot 0.361 79.56 18.24 2.20 23.72 24.03 33.67 37.10
Qwen3-32B baseline zero-shot 0.486 70.31 20.72 8.97 24.68 35.06 60.88 21.61
1-shot 0.418 74.95 18.60 6.45 23.72 27.60 53.06 25.81
5-shot 0.341 81.40 15.58 3.02 20.51 23.38 40.14 30.97
10-shot 0.309 82.01 15.17 2.83 19.23 21.75 32.99 30.65
Table 26: Error rates, incorrect choice distributions, and local negation confusion rates for the Qwen3 family under zero-shot, few-shot, and SFT conditions, evaluated in the completion-based setting using detailed instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
@@
Qwen3-0.6B-Base baseline zero-shot 0.658 48.80 4.46 46.75 27.88 37.66 38.78 28.39
1-shot 0.606 53.14 7.46 39.40 25.00 36.04 47.62 24.84
5-shot 0.588 52.77 3.24 43.99 26.60 33.12 43.20 25.48
10-shot 0.537 61.00 3.55 35.45 30.45 33.77 46.60 24.84
after SFT zero-shot 0.675 37.25 30.79 31.96 23.08 27.92 31.63 21.29
Qwen3-0.6B baseline zero-shot 0.541 74.93 7.92 17.16 43.91 40.91 57.48 25.48
1-shot 0.588 53.37 15.36 31.27 25.96 29.87 45.58 28.71
5-shot 0.535 65.04 6.81 28.15 31.09 31.49 52.72 29.03
10-shot 0.502 65.72 6.64 27.65 31.73 29.87 52.72 22.58
after SFT zero-shot 0.474 64.55 20.57 14.88 27.56 29.22 46.60 23.55
Qwen3-1.7B-Base baseline zero-shot 0.517 57.21 13.50 29.29 26.92 28.25 49.32 18.39
1-shot 0.457 69.97 13.89 16.15 31.73 32.79 40.14 27.42
5-shot 0.382 71.78 12.86 15.35 29.49 25.97 38.10 20.00
10-shot 0.343 72.29 11.55 16.17 25.32 23.70 36.73 17.10
after SFT zero-shot 0.391 68.15 13.18 18.66 30.45 26.30 41.16 12.58
Qwen3-1.7B baseline zero-shot 0.553 71.74 18.94 9.33 37.50 37.99 53.74 34.84
1-shot 0.508 68.12 14.06 17.81 30.45 31.82 55.44 25.81
5-shot 0.450 72.31 10.41 17.28 27.56 25.97 56.80 24.84
10-shot 0.455 71.25 10.10 18.64 30.13 25.65 55.78 23.23
after SFT zero-shot 0.427 66.05 26.90 7.05 29.81 29.87 34.35 22.58
Qwen3-4B-Base baseline zero-shot 0.416 79.01 14.89 6.11 36.86 24.68 48.64 25.81
1-shot 0.413 76.20 18.23 5.57 41.35 27.92 42.86 18.06
5-shot 0.366 75.49 17.57 6.94 39.42 25.97 30.27 18.06
10-shot 0.355 76.06 17.00 6.94 40.71 26.30 27.21 16.77
after SFT zero-shot 0.290 84.38 11.78 3.84 28.85 15.91 39.80 16.77
Qwen3-4B baseline zero-shot 0.424 81.27 17.23 1.50 37.18 29.22 51.36 24.84
1-shot 0.364 79.52 18.52 1.96 35.26 24.03 41.16 19.35
5-shot 0.347 82.88 14.61 2.51 32.37 26.30 38.10 22.26
10-shot 0.317 82.00 14.50 3.50 33.97 23.70 29.93 19.68
after SFT zero-shot 0.289 81.32 17.58 1.10 28.21 21.43 35.37 12.26
Qwen3-8B-Base baseline zero-shot 0.332 78.04 12.41 9.55 29.81 24.68 36.05 16.77
1-shot 0.305 77.86 13.80 8.33 30.45 26.62 25.51 15.16
5-shot 0.258 75.08 16.31 8.62 27.24 20.13 18.37 13.87
10-shot 0.251 78.23 12.93 8.83 27.24 19.81 18.03 15.81
after SFT zero-shot 0.177 74.89 16.14 8.97 17.63 14.61 9.52 12.58
Qwen3-8B baseline zero-shot 0.381 86.07 12.27 1.66 43.59 33.12 37.41 21.29
1-shot 0.305 83.38 14.03 2.60 36.54 24.03 26.87 17.42
5-shot 0.297 81.87 13.60 4.53 34.62 26.30 21.77 17.42
10-shot 0.276 81.32 12.93 5.75 31.73 23.70 18.71 18.06
after SFT zero-shot 0.240 85.48 10.56 3.96 29.17 21.43 19.05 14.84
Qwen3-14B-Base baseline zero-shot 0.264 79.58 15.92 4.50 20.83 16.56 27.89 21.61
1-shot 0.242 73.11 21.97 4.92 21.79 16.88 20.75 13.55
5-shot 0.210 75.09 21.51 3.40 22.12 15.58 13.95 13.23
10-shot 0.182 77.29 20.52 2.18 18.59 15.58 8.50 14.84
Qwen3-14B baseline zero-shot 0.280 79.04 19.55 1.42 28.85 23.05 25.17 14.19
1-shot 0.291 80.65 15.53 3.81 31.73 21.10 29.93 14.19
5-shot 0.241 86.51 11.18 2.30 30.13 20.78 20.75 14.19
10-shot 0.241 84.21 13.82 1.97 31.73 21.75 16.33 13.55
Qwen3-32B baseline zero-shot 0.240 78.81 19.54 1.66 20.83 24.35 18.03 14.52
1-shot 0.189 80.67 18.49 0.84 22.44 20.45 10.88 8.71
5-shot 0.194 80.82 18.78 0.41 21.47 20.45 12.93 9.68
10-shot 0.175 80.54 19.00 0.45 19.23 18.18 9.18 11.29
Table 27: Error rates, incorrect choice distributions, and local negation confusion rates for the Qwen3 family under zero-shot, few-shot, and SFT conditions, evaluated in the option-selection setting using detailed instruction.
Model Training Setting N Shot Error Rate (1-acc) Incorrect Choice Distribution Local Negation Confusion Rate
@@
Llama-3.1-8B baseline zero-shot 0.536 76.92 17.60 5.47 25.32 30.19 60.20 55.16
1-shot 0.501 80.06 16.14 3.80 24.68 29.22 57.14 55.16
5-shot 0.391 83.77 14.20 2.03 21.79 23.05 43.54 47.10
10-shot 0.356 83.07 15.14 1.78 19.23 21.43 38.78 42.90
after SFT zero-shot 0.226 88.07 10.18 1.75 10.58 15.91 24.15 31.61
Llama-3.1-8B- Instruct baseline zero-shot 0.536 73.82 23.96 2.22 25.00 31.82 58.50 48.71
1-shot 0.457 79.51 19.10 1.39 24.36 28.57 48.98 48.39
5-shot 0.367 85.31 14.25 0.43 18.91 22.40 40.14 48.06
10-shot 0.332 85.17 14.35 0.48 20.51 20.78 32.99 42.26
after SFT zero-shot 0.229 86.51 12.46 1.04 10.90 14.29 23.13 33.55
Llama-3.2-1B baseline zero-shot 0.591 71.01 24.03 4.97 21.79 29.55 63.95 58.71
1-shot 0.570 72.32 23.37 4.31 25.32 32.47 63.95 49.35
5-shot 0.514 75.77 20.52 3.70 20.83 27.27 63.95 49.68
10-shot 0.468 78.14 18.64 3.22 19.87 25.97 59.86 46.13
after SFT zero-shot 0.250 76.83 21.59 1.59 13.46 14.61 24.15 27.10
Llama-3.2-1B- Instruct baseline zero-shot 0.569 68.48 27.89 3.63 25.00 30.84 58.16 47.42
1-shot 0.526 69.83 27.75 2.41 25.32 26.95 54.76 45.16
5-shot 0.472 70.92 25.38 3.70 21.15 24.68 51.36 41.61
10-shot 0.447 72.70 23.94 3.37 22.12 22.40 47.96 42.26
after SFT zero-shot 0.239 81.06 17.94 1.00 13.14 15.26 21.43 30.00
Llama-3.2-3B baseline zero-shot 0.565 74.58 19.94 5.48 24.36 31.49 62.93 55.81
1-shot 0.538 78.17 18.88 2.95 31.09 31.82 60.88 50.32
5-shot 0.429 79.85 17.56 2.59 22.44 27.27 47.96 44.19
10-shot 0.413 81.96 15.16 2.88 23.72 26.95 45.92 43.55
after SFT zero-shot 0.244 84.36 14.33 1.30 10.58 16.88 24.83 32.58
Llama-3.2-3B- Instruct baseline zero-shot 0.548 76.70 21.71 1.59 27.88 33.12 54.42 58.39
1-shot 0.536 76.33 22.34 1.33 31.41 31.17 52.72 53.87
5-shot 0.448 78.76 20.18 1.06 24.68 26.95 40.82 53.23
10-shot 0.414 79.12 19.73 1.15 24.04 24.35 37.41 49.35
after SFT zero-shot 0.205 84.50 13.18 2.33 8.33 10.06 25.17 28.06